OpenStudy (anonymous):

Which of the following is equivalent to log(24)8?

OpenStudy (anonymous):

a. log(24)/log(8) b. log(8)/log(24) c. log 3 d. log 1/3

OpenStudy (anonymous):

change of base formula \[\large \log_b(x)=\frac{\log(x)}{\log(b)}\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

so b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ty